OpenStudy (shelbywyatt):

Suppose "p" varies directly with "d", and "p=3" and d=5. What is the value of "d" when "p=12"

OpenStudy (niksbhalla14mar):

D =20.

OpenStudy (niksbhalla14mar):

if d varies directly with d then p=k.d For p=3 and d=5. k =3/5=0.6 If p =12 and k =.6 Then in equation p=k.d , d= p/k =12/0.6 = 20

OpenStudy (shelbywyatt):

thanks